# Seat Arrangement Problems under B-utility and W-utility

### v1

@@ Abstract
In the Seat Arrangement problem the goal is to allocate agents to vertices in a graph such that the resulting arrangement is optimal or fair in some way. Examples include an arrangement that maximises utility or one where no agent envies another. We introduce two new ways of calculating the utility that each agent derives from a given arrangement, one in which agents care only about their most preferred neighbour under a given arrangement, and another in which they only care about their least preferred neighbour. We also present a new restriction on agent’s preferences, namely $1$-dimensional preferences. We give algorithms, hardness results, and impossibility results for these new types of utilities and agents’ preferences. Additionally, we refine previous complexity results, by showing that they hold in more restricted settings.
@@ 1 Introduction
Let us imagine that we are organizing a wedding. An important part of the event is the dinner, and for that we need to decide where our guests are seated. However, guests have preferences over who sits next to them. For example, couples want to be together, parents may prefer to be next to their kids, and kids may want to be next to other kids. We may even have the case of two relatives who dislike each other, and thus would rather not be seated together. The problem of deciding how to best allocate these wedding guests to seats is known as Seat Arrangement [4].
@@
Up until now, the only type of utility that had been considered is one where the utility of an agent equals the sum of the valuations that said agent gives to each of their neighbours, which we call S-utility. However, depending on the setting, agents may only care about who their best neighbour is, or conversely who their worst neighbour is. In this work we consider these two types of utilities, referred to as B-utility and W-utility respectively, that are new concepts in the context of Seat Arrangement.
Additionally, because most problems related to Seat Arrangement are computationally hard, we present a new restriction on the preferences of agents, referred to as $1$-dimensional preferences. Under this restriction, agents’ preferences must be derived from a one-dimensional space. An example would be political ideology, where left-wing agents prefer other left-wing agents over right-wing agents, and similarly for agents located at other points in the left-right axis. Other restrictions on preferences that we study are strict preferences, where no agent can give the same valuation to two different agents; binary preferences, in which the only allowed valuations are zero and one; and symmetric preferences, where the valuation that an agent $p$ gives to an agent $q$ is equal to the valuation that $q$ gives to $p$.
@@ 1.1 Our contribution
In this paper we introduce two different ways of calculating the utility that an agent derives from a seating arrangement: B-utility and W-utility. We study the problems of finding optimal arrangements in instances that have one of these two utilities, and present both algorithms and computational hardness results. In particular, we show that with B-utility and W-utility, MWA, MUA, and EFA are NP-hard, even if preferences are symmetric or binary, or symmetric and strict. On the other hand, if preferences are symmetric, we prove that with either B-utility or W-utility an exchange-stable arrangement is guaranteed to exist, although only in the first case the algorithm is shown to run in polynomial time.
Seat Arrangement with the most common utility, S-utility, is also studied, and many previous complexity results are refined. Thus, we prove that, with S-utility, MWA, MUA and EFA are NP-hard even if preferences are symmetric and strict. EFA remains NP-hard if preferences are symmetric and binary. Another contribution is the introduction of a type of agents’ preferences, $1$-dimensional preferences, new in the Seat Arrangement problem. We show that if the seat graph has a connected component which is a cycle with at least four vertices, or a path with at least three vertices, then no envy-free arrangement exists, if the utility type of the instance is S-utility. If the utility type is B-utility, and preferences are strict and $1$-dimensional, MWA, MUA, and EFA are NP-hard.
@@
MWA MUA EFA STA Pref. Graph B S W B S W B S W B S W Arb. N [T 9] N [4] N [T 9] N [T 9] N [4] N [T 9] N [T 3] N[16] N [T 5] N [O 1] N [4, 6, 9] N [O 1] Arb. Match. P[O 1] P [4] P[O 1] P [O 1] P [4] P [O 1] P [O 1] P [4] P [O 1] N [O 1] N [4, 6, 9] N [O 1] Strict Arb. N [T 9] N [T 9] N [T 9] N [T 9] N [T 9] N [T 9] N [T 8] N [T 4] N [T 6] N [O 1] N [4, 6, 9] N [O 1] Symm. Arb. N [T 9] N [4] N [T 9] N [T 9] N [4][T 9] N [T 9] N [T 3] N [4] N [T 5] P, $\exists$ [T 2] $\exists$ [16] $\exists$ [T 1] Binary Arb. N [T 9] N [4] N [T 9] N [T 9] N [4][T 9] N [T 9] N [T 3] N [16] N [T 5] Paths N [P 2] N [P 2] N [T 8] $\nexists$ [T 7] P, $\exists$ [T 2] P, $\exists$ [P1] P, $\exists$ [P1] 1-D Cycles N [P 2] N [P 2] N [T 8] $\nexists$ [T 7] P, $\exists$ [T 2] P, $\exists$ [P1] P, $\exists$ [P1]
Table 1: Summary of results for Seat Arrangement. Under each of the four problems (MWA, MUA,…), we have results for each of the three utility types. B stands for B-utility, S for S-utility and W for W-utility. Similarly, in the first column, Pref stands for Preferences, Arb means arbitrary preferences, and Symm means symmetric preferences. In the second column Match means matching graph. Cells with N indicate that the problem is NP-hard, while cells with P indicate that the problem can be solved in polynomial time. In the case of EFA and STA, if a type of arrangement is guaranteed to exist, its cell in the table will contain $\exists$. If the arrangement is guaranteed not to exist, its cell will contain $\nexists$. The results included in this paper are preceded by a letter, which indicates if the result is a theorem (T), an observation (O), or a proposition (P)
@@ 1.2 Related work
The Seat Arrangement problem was first considered by Bodlaender et al. [4]. However, similar problems, such as exchange-stable roommates [1, 6, 9], had already been studied prior to that work. In an exchange-stable roommates instance there is a set of agents, and the goal is to partition the set into subsets of size two (pairs) such that no two agents want to swap their partners, in which case we have an exchange-stable matching. An application is the allocation of housing in college, in which each room houses two students. A matching is exchange-stable if no two students assigned to different rooms wish to switch their rooms, and hence their roommates. Thus, this problem is equivalent to determining the existence of an exchange-stable arrangement in Seat Arrangement, if we restrict ourselves to seat graphs where every connected component has two vertices. An important result in exchange-stable roommates is that an exchange-stable matching need not exist and deciding whether one does exist is NP-complete [6, 9]. As Seat Arrangement is a generalization of exchange-stable roommates, this result holds for Seat Arrangement too. A related problem is Stable Roommates [13], in which the goal is to partition a set of agents into pairs, and a pair of agents blocks if they wish to be with each other instead of their respective partners in the matching. In contrast to the case for exchange-stable roommates, we can find a stable matching, i.e., a matching with no blocking pairs, or report that none exists, in polynomial time [15].
@@
Another important result for Seat Arrangement comes from Massand and Simon [16]. In their paper they study a generalization of Seat Arrangement where agents derive an utility not just from the sum of their valuations over their neighbours, but also from an intrinsic valuation of the seat assigned to them. They show that, if preferences are symmetric, an exchange-stable arrangement always exists, and can be found by successively switching pairs of agents that desire to exchange seats, although this algorithm is PLS-complete. However, with arbitrary preferences, it was shown that finding an envy-free arrangement is NP-hard, even in the case where the valuation of every vertex is the same, meaning that EFA is hard in Seat Arrangement too.
@@
Bodlaender et al. [4] considered Seat Arrangement with S-utility, and showed that, for EFA, hardness remains even if preferences are symmetric and every graph component has at most three vertices. Likewise, MWA, the problem of finding an arrangement that maximises the total utility, is NP-hard, even for symmetric and binary preferences, for a large class of graphs. The same result applies to MUA, the problem of finding an arrangement that maximises the utility of the agent with the lowest utility, but only if Spanning Subgraph Isomorphism is NP-complete in regular graphs, which is currently unknown. On the other hand, if every component of the seat graph has two vertices, then EFA, MWA and MUA are solvable in polynomial time.
@@
Due to the computational hardness of most problems related to Seat Arrangement, there has been some work on parameterized complexity, and in finding fixed-parameter tractable algorithms in particular [12]. For instance, Bodlaender et al. [4] studied the parameterized complexity of MWA, MUA, EFA, and STA where the parameter is the vertex cover number of the seat graph. They show, among other results, that MWA parameterized by the vertex cover number is W[1]-hard, and that both MUA and EFA are weakly NP-hard even if the vertex cover number is two. Ceylan et al. also gave parameterized complexity results for Seat Arrangement with S-utility [11]. They considered two parameters, the number of non-isolated vertices in the seat graph, and the maximum number of non-zero valuations that any agent can give to other agents, and provide both FPT algorithms [12] and W[1]-hardness [12] results. Another approach, which we also take on this work, is to restrict the preferences of agents. For instance, Berriaud et al. [3] study envy-freeness and exchange stability in instances with path and cycle graphs, where the agents can be partitioned into a number of classes, such that agents within one class are equivalent to each other. They find that these problems are solvable in polynomial time if the number of classes is bounded, but they remain NP-hard if they are not [3].
@@
Up until now, in all papers related to Seat Arrangement, the only type of agent utility considered was S-utility. However, that is not the case in the related problem of coalition formation with hedonic preferences [2], where the goal is to find a partition of a given set of agents that satisfies a concept of stability that depends on the agents’ preferences. In such problems, the concept of agents only caring about their preferred, or least-preferred, agent in the partition is present in much of the literature [10, 7, 8]. Determining the existence of such partitions is usually possible in polynomial time, but some variants are NP-hard. Coalition formation with hedonic preferences is related to Seat Arrangement, as an instance of the latter in which every connected component of the seat graph is a clique can be understood as an instance of coalition formation with hedonic preferences.
@@ 2 Preliminaries
A Seat Arrangement instance $I=(P,F,G,U)$ has four elements: a set $P$ of agents, a set $F$ of valuation functions $f_{p}:P\backslash\{p\}\rightarrow\mathbb{R}$ for each $p\in P$ that determines $p$’s preferences over all other agents, an undirected graph $G$, with as many vertices as there are agents and where the agents will be placed, called the seat graph, and a utility type $U$, which indicates the mechanism that is used to derive a utility for an agent depending on their placement in the seat graph. Our objective is to find an assignment from agents to vertices of $G$ that meets an optimality or fairness goal.
Let $p,q,r\in P$, $p\notin\{q,r\}$. The valuation that $p$ gives to $q$ is the number $f_{p}(q)$. We say that $p$ strictly prefers $q$ over $r$ if $f_{p}(q)>f_{p}(r)$, and is indifferent between $q$ and $r$ if $f_{p}(q)=f_{p}(r)$.
@@
  • $1$-dimensional if agents are points in the real line. Here, $d(p,q)$ represents the distance between $p$ and $q$ in the line, and $D$ denotes the distance between the leftmost and the rightmost agent. Then $\forall p\neq q\in P$, $f_{p}(q)=D-d(p,q)+1$. If no two agents are at the same point in the real line, their positions are unique. $1$-dimensional preferences are, by definition, positive and symmetric.
@@
We define an arrangement (alternatively an assignment) to be a bijection $\sigma:P\rightarrow V(G)$. Thus $\sigma(p)\in G$ is the vertex in $G$ that agent $p$ is assigned to. Alternatively, $\sigma(p)$ is the seat of $p$ under $\sigma$. If agents’ preferences are symmetric, we say that an edge $e\in E(G)$ has score $s\in\mathbb{R}$ under $\sigma$ if $e=\{\sigma(p),\sigma(q)\}$ for $p,q\in P$, and the pair $\{p,q\}$ has score $s$. For any arrangement $\sigma$ and agents $p,q\in P$, $p\neq q$, let $\sigma^{p\leftrightarrow q}$ be the arrangement where
@@
A function $u_{\sigma}:P\rightarrow\mathbb{R}$ assigns a real number to every agent, which represents the utility that the agent derives under arrangement $\sigma$. We will consider three different types of utility functions. We say that $u_{\sigma}$ is an additively separable utility function if $u_{\sigma}(p)=\sum_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))$, that is, the utility of $p$ under $\sigma$ is the sum of the valuations under $f_{p}$ for each of the agents neighbouring $p$. We will refer to additively separable utility functions as S-utility functions also (S stands for sum). Alternatively, $u_{\sigma}$ is a B-utility (respectively W-utility) function if $u_{\sigma}(p)=\max_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))$ (resp. $\min_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))$). In other words, under B-utility, agent $p$ cares only about its best neighbour, while under W-utility, $p$ cares only about its worst neighbour. In a Seat Arrangement instance $I=(P,F,G,U)$, $U$ specifies the type of all utility functions in the instance, with $U\in\{B,S,W\}$.
We will now consider solution concepts corresponding to optimal or fair arrangements. First, we define two specific types of fair arrangement. Under arrangement $\sigma$, we say that agent $p$ envies agent $q$ if $u_{\sigma}(p)<u_{\sigma^{p\leftrightarrow q}}(p)$. Agents $p$ and $q$ form an exchange-blocking pair if they envy each other. An arrangement $\sigma$ is called envy-free if no agent $p$ envies another agent $q$, and $\sigma$ is called exchange-stable if no exchange-blocking pair exists. Envy-freeness implies exchange-stability.
@@
In a matching graph, the degree of any vertex is $1$. Thus, for a given agent $p$ and arrangement $\sigma$, $\max_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))=\sum_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))=\min_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))$, i.e., $p$ has the same utility under all three different utilities because there is only one neighbour. ∎
Therefore, for both B- and W-utility, finding an exchange-stable arrangement is NP-hard if preferences are arbitrary [4, 6, 9], while if the seat graph is a matching graph, EFA, MWA and MUA are solvable in polynomial time [4].

### v2

@@ Abstract
In the seat arrangement problem the goal is to allocate agents to vertices in a graph such that the resulting arrangement is optimal or fair in some way. Examples include an arrangement that maximises utility or one where no agent envies another. We introduce two new ways of calculating the utility that each agent derives from a given arrangement, one in which agents care only about their most preferred neighbour under a given arrangement, and another in which they only care about their least preferred neighbour. We also present a restriction on agent’s preferences, namely $1$-dimensional preferences. We give algorithms, hardness results, and impossibility results for these types of utilities and agents’ preferences. Additionally, we refine previous complexity results, by showing that they hold in more restricted settings.
@@ 1 Introduction
Let us imagine that we are organizing a wedding. An important part of the event is the dinner, and for that we need to decide where our guests are seated. However, guests have preferences over who sits next to them. For example, couples want to be together, parents may prefer to be next to their kids, and kids may want to be next to other kids. We may even have the case of two relatives who dislike each other, and thus would rather not be seated together. The problem of deciding how to best allocate these wedding guests to seats is known as seat arrangement [bodlaender_hedonic_2020].
@@
Up until now, the only type of utility that had been considered is one where the utility of an agent equals the sum of the valuations that said agent gives to each of their neighbours, which we call S-utility. However, depending on the setting, agents may only care about who their best neighbour is, or conversely who their worst neighbour is. For example, in a wedding, guests may be happy being seated next to only one person whom they like, as each guest could spend the dinner talking to their favourite neighbour and ignore the others. On the other hand, in an office setting, employers may be keen to avoid sitting two employees who dislike each other together, even if each of them is surrounded by coworkers whom they like, as these two problematic employees may be unhappy being next to each other, and their work could suffer. In this work we consider these two types of utilities, referred to as B-utility and W-utility respectively, that are new concepts in the context of seat arrangement.
Additionally, because most problems related to seat arrangement are computationally hard, we present a restriction on the preferences of agents, referred to as $1$-dimensional preferences. Under this restriction, agents’ preferences must be derived from a $1$-dimensional space. An example would be political ideology, where left-wing agents prefer other left-wing agents over right-wing agents, and similarly for agents located at other points in the left-right axis. Other restrictions on preferences that we study are strict preferences, where no agent can give the same valuation to two different agents; binary preferences, in which the only allowed valuations are zero and one; and symmetric preferences, where the valuation that an agent $p$ gives to an agent $q$ is equal to the valuation that $q$ gives to $p$.
@@ 1.1 Our contribution
In this paper we introduce two different ways of calculating the utility that an agent derives from a seating arrangement: B-utility and W-utility. We study the problems of finding optimal arrangements in instances that have one of these two utilities, and present both algorithms and computational hardness results. In particular, we show that with B-utility and W-utility, MWA, MUA, and EFA are NP-hard, even if preferences are symmetric and binary, or symmetric and strict. On the other hand, if preferences are symmetric, we prove that with both B-utility or W-utility an exchange-stable arrangement is guaranteed to exist, although only in the first case the algorithm is shown to run in polynomial time.
seat arrangement with the most common utility, S-utility, is also studied, and many previous complexity results are refined. Thus, we prove that, with S-utility, MWA, MUA and EFA are NP-hard even if preferences are symmetric and strict. Another contribution is the introduction of a type of agents’ preferences, $1$-dimensional preferences. We show that if the seat graph has a connected component which is a cycle with at least four vertices, or a path with at least three vertices, then no envy-free arrangement exists, if the utility type of the instance is S-utility. If the utility type is B-utility, and preferences are strict and $1$-dimensional, MWA, MUA, and EFA are NP-hard.
@@
MWA MUA EFA STA Val. Graph B S W B S W B S W B S W Arb. N [T 9] N [bodlaender_hedonic_2020] N [T 9] N [T 9] N [bodlaender_hedonic_2020] N [T 9] N [T 3] N[massand_graphical_2019] N [T 6] N [O 1] N [bodlaender_hedonic_2020, cechlarova_complexity_2002, cechlarova_exchange-stable_2005] N [O 1] Arb. Match. P[O 1] P [bodlaender_hedonic_2020] P[O 1] P [O 1] P [bodlaender_hedonic_2020] P [O 1] P [O 1] P [bodlaender_hedonic_2020] P [O 1] N [O 1] N [bodlaender_hedonic_2020, cechlarova_complexity_2002, cechlarova_exchange-stable_2005] N [O 1] Strict Arb. N [T 9] N [T 9] N [T 9] N [T 9] N [T 9] N [T 9] N [T 4] N [T 5] N [T 7] N [O 1] N [bodlaender_hedonic_2020, cechlarova_complexity_2002, cechlarova_exchange-stable_2005] N [O 1] Symm. Arb. N [T 9] N [bodlaender_hedonic_2020] N [T 9] N [T 9] N [bodlaender_hedonic_2020][T 9] N [T 9] N [T 3] N [bodlaender_hedonic_2020, ceylan_optimal_2022] N [T 6] P, $\exists$ [T 2] $\exists$ [massand_graphical_2019] $\exists$ [T 1] Binary Arb. N [T 9] N [bodlaender_hedonic_2020] N [T 9] N [T 9] N [bodlaender_hedonic_2020][T 9] N [T 9] N [T 3] N [massand_graphical_2019, ceylan_optimal_2022] N [T 6] Open Open Open Paths N [P 2] Open Open N [P 2] Open Open N [T 4] $\nexists$ [T 8] Open P, $\exists$ [T 2] P, $\exists$ [P1] P, $\exists$ [P1] 1-D Cycles N [P 2] Open Open N [P 2] Open Open N [T 4] $\nexists$ [T 8] Open P, $\exists$ [T 2] P, $\exists$ [P1] P, $\exists$ [P1]
Table 1: Summary of results for seat arrangement. Under each of the four problems (MWA, MUA,…), we have results for each of the three utility types. B stands for B-utility, S for S-utility and W for W-utility. Similarly, in the first column, Val stands for valuations, Arb means arbitrary seat graph, and Symm means symmetric valuations. In the second column Match means matching graph. Cells with N indicate that the problem is NP-hard, while cells with P indicate that the problem can be solved in polynomial time. In the case of EFA and STA, if a type of arrangement is guaranteed to exist, its cell in the table will contain $\exists$. If the arrangement never exists, its cell will contain $\nexists$. The results included in this paper are preceded by a letter, which indicates if the result is a theorem (T), an observation (O), or a proposition (P).
@@ 1.2 Related work
The seat arrangement problem was first considered by Bodlaender et al. [bodlaender_hedonic_2020]. However, similar problems, such as exchange-stable roommates [alcalde_exchange-proofness_1994, cechlarova_complexity_2002, cechlarova_exchange-stable_2005], had already been studied prior to that work. In an exchange-stable roommates instance there is a set of agents, and the goal is to partition the set into subsets of size two (pairs) such that no two agents want to swap their partners, in which case we have an exchange-stable matching. An application is the allocation of housing in college, in which each room houses two students. A matching is exchange-stable if no two students assigned to different rooms wish to switch their rooms, and hence their roommates. Thus, this problem is equivalent to determining the existence of an exchange-stable arrangement in seat arrangement, if we restrict ourselves to seat graphs where every connected component has two vertices. An important result in exchange-stable roommates is that an exchange-stable matching need not exist and deciding whether one does exist is NP-complete [cechlarova_complexity_2002, cechlarova_exchange-stable_2005]. As seat arrangement is a generalization of exchange-stable roommates, this result holds for seat arrangement too. A related problem is stable roommates [gale_college_1962], in which the goal is to partition a set of agents into pairs, and a pair of agents blocks if they wish to be with each other instead of their respective partners in the matching. In contrast to the case for exchange-stable roommates, we can find a stable matching, i.e., a matching with no blocking pairs, or report that none exists, in polynomial time [gusfield_stable_1989]. Some other related problems include hedonic games [brandt_handbook_2016], Schelling games [elkind_2019_schelling], and topological distance games [bullinger_2024_topological].
@@
Another important result for seat arrangement comes from Massand and Simon [massand_graphical_2019]. In their paper they study a generalization of seat arrangement where agents derive an utility not just from the sum of their valuations over their neighbours, but also from an intrinsic valuation of the seat assigned to them. They show that, if preferences are symmetric, an exchange-stable arrangement always exists, and can be found by successively switching pairs of agents that desire to exchange seats, although the problem is PLS-complete. However, with arbitrary preferences, it was shown that finding an envy-free arrangement is NP-hard, even in the case where the valuation of every vertex is the same, meaning that EFA is hard in seat arrangement too.
@@
Bodlaender et al. [bodlaender_hedonic_2020] considered seat arrangement with S-utility, and showed that, for EFA, hardness remains even if preferences are symmetric and every graph component has at most three vertices. Likewise, MWA, the problem of finding an arrangement that maximises the total utility, is NP-hard, even for symmetric and binary preferences, for a large class of graphs. The same result applies to MUA, the problem of finding an arrangement that maximises the utility of the agent with the lowest utility, but only if Spanning Subgraph Isomorphism is NP-complete in regular graphs, which is currently unknown. On the other hand, if every component of the seat graph has two vertices, then EFA, MWA and MUA are solvable in polynomial time.
@@
Due to the computational hardness of most problems related to seat arrangement, there has been some work on parameterized complexity, and in finding fixed-parameter tractable algorithms in particular [cygan_parameterized_2015]. For instance, Bodlaender et al. [bodlaender_hedonic_2020] studied the parameterized complexity of MWA, MUA, EFA, and STA where the parameter is the vertex cover number of the seat graph. They show, among other results, that MWA parameterized by the vertex cover number is W[1]-hard, and that both MUA and EFA are weakly NP-hard even if the vertex cover number is two. Ceylan et al. also gave parameterized complexity results for seat arrangement with S-utility [ceylan_optimal_2023]. They considered two parameters, the number of non-isolated vertices in the seat graph, and the maximum number of non-zero valuations that any agent can give to other agents, and provide both FPT algorithms and W[1]-hardness results.
Another approach, which we also take on this work, is to restrict the preferences of agents. For instance, Berriaud et al. [berriaud_stable_2023] study envy-freeness and exchange stability in instances with path and cycle graphs, where the agents can be partitioned into a number of classes, such that agents within one class are equivalent to each other. They find that these problems are solvable in polynomial time if the number of classes is bounded, but they remain NP-hard if they are not [berriaud_stable_2023]. Similarly, Ceylan studied envy-freeness, and showed that EFA is NP-hard under S-utility, even with symmetric and binary preferences [ceylan_optimal_2022].
Another restriction of agents’ preferences is $1$-dimensional preferences. However, with the exception of a paper by Wilczynski [wilczynski_ordinal_2023], $1$-dimensional preferences had not been studied before in the context of seat arrangement. Moreover, in that paper, $1$-dimensional preferences were only considered in the problem of finding popular arrangements, so no previous results for $1$-dimensional preferences and STA, EFA, MWA or MUA are known [^1] Positive results with arbitrary or symmetric preferences extend to $1$-dimensional preferences.. Nevertheless, $1$-dimensional preferences, and generalizations like single-peaked preferences or $d$-dimensional preferences (for $d\geq 1$), have been studied before in different matching and social choice problems. Some examples are stable marriage and stable roommates [arkin_geometric_2009, bartholdi_stable_1986, bredereck_stable_2020, chen_et_al:LIPIcs.ESA.2022.36, wen_stable_2023], voting theory [brandt_handbook_2016, black1948rationale, brandt_2010_bypassing], and resource allocation problems [sprumont_1991_division].
@@
Up until now, in all papers related to seat arrangement, the only type of agent utility considered was S-utility. However, that is not the case in the related problem of coalition formation with hedonic preferences [brandt_handbook_2016], where the goal is to find a partition of a given set of agents that satisfies a concept of stability that depends on the agents’ preferences. In such problems, the concept of agents only caring about their preferred, or least-preferred, agent in the partition is present in much of the literature [cechlarova_computational_2003, cechlarova_stable_2004, cechlarova_stability_2001]. Determining the existence and finding such partitions is usually possible in polynomial time, but some variants are NP-hard. Coalition formation with hedonic preferences is related to seat arrangement, as an instance of the latter in which every connected component of the seat graph is a clique can be understood as an instance of coalition formation with hedonic preferences, as every connected component would be a distinct coalition. Nevertheless, results, positive or otherwise, from one problem do not translate to the other. The main difference is that, in coalition formation with hedonic preferences, a coalition of agents blocks  [^2] Other stability concepts do exist, see here [brandt_handbook_2016] for an overview, but they are generally not equivalent to exchange-stability. if each agent can break away from their assigned coalition and form a new coalition in which they are better off. However, this is not possible in seat arrangement, as an agent can only get a new seat if they swap seats with another agent. This point is also what differentiates stable roommates from exchange-stable roommates.
@@ 2 Preliminaries
[…]
Let $\in P, p{q,r}p\notin\{q,r\}$. The valuation that $p$ gives to $q$ is the number $f_{p}(q)$. We say that $p$ strictly prefers $q$ over $r$ if $f_{p}(q)>f_{p}(r)$, and is indifferent between $q$ and $r$ if $f_{p}(q)=f_{p}(r)$.
@@
  • 11-dimensional if agents are points in the real line. Here, $d(p,q)$ represents the distance between $p$ and $q$ in the line, and $D$ denotes the distance between the leftmost and the rightmost agent. Then $\forall p\neq q\in P$, $f_{p}(q)=D-d(p,q)+1$. $1$-dimensional preferences are, by definition, positive, because $D\geq d(p,q)$, and the +1+1 term ensures that $f_{p}(q)>0$. Likewise, they are symmetric, as $f_{p}(q)=D-d(p,q)+1=D-d(q,p)+1=f_{q}(p)$. If no two agents are at the same point in the real line, their positions are unique. We refer to the position or location of agent $a$ in the $1$-dimensional space as $l_{a}$, so if $l_{a}=1$ then $a$ is at position $1$.
@@
We define an arrangement (alternatively an assignment) to be a bijection $\sigma:P\rightarrow V(G)$. Thus $\sigma(p)\in G$ is the vertex in $G$ that agent $p$ is assigned to. Alternatively, $\sigma(p)$ is the seat of $p$ under $\sigma$. If agents’ preferences are symmetric, we say that an edge $e\in E(G)$ has score $s\in\mathbb{R}$ under $\sigma$ if $e=\{\sigma(p),\sigma(q)\}$ for $p,q\in P$, and the pair $\{p,q\}$ has score $s$. For any arrangement $\sigma$ and agents $\in P, pqp\neq q$, let $\sigma^{p\leftrightarrow q}$ be the arrangement where
@@
A function $u_{\sigma}:P\rightarrow\mathbb{R}$ assigns a real number to every agent, which represents the utility that the agent derives under arrangement $\sigma$. We will consider three different types of utility functions. We say that $u_{\sigma}$ is an additively separable utility function if $u_{\sigma}(p)=\sum_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))$, where $N(\sigma(p))$ is the set that includes all vertices adjacent to $\sigma(p)$. That is, the utility of $p$ under $\sigma$ is the sum of the valuations under $f_{p}$ for each of the agents neighbouring $p$. We will refer to additively separable utility functions as S-utility functions also (S stands for sum). Alternatively, $u_{\sigma}$ is a B-utility (respectively W-utility) function if $u_{\sigma}(p)=\max_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))$ (resp. $\min_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))$). In other words, under B-utility, agent $p$ cares only about its best neighbour, while under W-utility, $p$ cares only about its worst neighbour. If an agent $p$ is assigned to a vertex with no neighbours, then its utility is 0 under all utility types. In a seat arrangement instance $I=(P,F,G,U)$, $U$ specifies the type of all utility functions in the instance, with $U\in\{B,S,W\}$.
We will now consider solution concepts corresponding to optimal or fair arrangements. First, we define two specific types of fair arrangement. Under arrangement $\sigma$, we say that agent $p$ envies agent $q$ if $u_{\sigma}(p)<u_{\sigma^{p\leftrightarrow q}}(p)$. Informally, $p$ envies $q$ if under arrangement $\sigma$ $p$ prefers to be seated where $q$ is seated. Agents $p$ and $q$ form an exchange-blocking pair if they envy each other. An arrangement $\sigma$ is called envy-free if no agent $p$ envies another agent $q$, and $\sigma$ is called exchange-stable if no exchange-blocking pair exists. Envy-freeness implies exchange-stability.
@@
In a matching graph, the degree of any vertex is $1$. Thus, for a given agent $p$ and arrangement $\sigma$, $\max_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))=\sum_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))=\min_{v\in N(\sigma(p))}f_{p}(\sigma^{-1}(v))$, i.e., $p$ has the same utility under all three different utility types because there is only one neighbour. ∎
Therefore, for both B-utility and W-utility, finding an exchange-stable arrangement is NP-hard if preferences are strict but otherwise arbitrary, even if the seat graph is a matching graph, because under S-utility finding a exchange-stable arrangement is NP-hard if preferences are strict and the seat graph is a matching graph [bodlaender_hedonic_2020, cechlarova_complexity_2002, cechlarova_exchange-stable_2005]. However, if the seat graph is a matching graph, EFA, MWA and MUA are solvable in polynomial time [bodlaender_hedonic_2020].